# On the Training Convergence of Transformers for In-Context Classification

### v1

@@ 4 In-context multi-class classification
$E=E(P)=\begin{pmatrix}x_{1}&x_{2}&\cdots&x_{N}&x_{\mathsf{query}}\\ y_{1}&y_{2}&\cdots&y_{N}&0_{c}\end{pmatrix}\in\mathbb{R}^{(d+c)\times(N+1)}.$ (13)
@@
$\widehat{y}_{\mathsf{out}}=\mathrm{softmax}\left([F(E(P);W)]_{(d+1):(d+c),(N+1)}\right)=\mathrm{softmax}\left(\left(\frac{1}{N}\sum_{i=1}^{N}y_{i}x_{i}^{\top}\right)Wx_{\mathsf{query}}\right).$
@@ 4.1 Training procedure
Assumption 4.1
For each learning task $\tau\in[B]$, we assume
@@
    $\{x_{\tau,i},y_{\tau,i}\}_{i=1}^{N},\{x_{\tau,\mathsf{query}},y_{\tau,\mathsf{query}}\}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{P}^{m}(\mu_{\tau}=(\mu_{\tau,1},\dots,\mu_{\tau,c}),\Lambda)$.
@@
    $\mu_{\tau,1}$ is sampled from $\mathsf{N}(0,I_{d})$, and $\mu_{\tau,k}=U_{\tau,k,\Lambda}\mu_{\tau,1},k=2,3,\dots,c$, where $U_{\tau,k,\Lambda}=\Lambda^{1/2}U_{\tau,k}\Lambda^{-1/2}$, and $U_{\tau,k}$ are uniformly distributed over the closed set of real unitary matrices such that $U_{\tau,k}U_{\tau,k}^{\top}=I_{d}$.
We denote the distribution of $\mu_{\tau}$ as $\mathcal{P}^{m}_{\Omega}(\Lambda)$. Note that $U_{\tau,k,\Lambda}=\Lambda^{1/2}U_{\tau,k}\Lambda^{-1/2}$ can be viewed as linear transformations that preserve the inner product of vectors in the $\Lambda^{-1}$ weighted norm, and we have $\mu_{\tau,i}^{\top}\Lambda^{-1}\mu_{\tau,i}=\mu_{\tau,j}^{\top}\Lambda^{-1}\mu_{\tau,j}$, for $i,j\in[c]$. Let $\widehat{y}_{\tau,\mathsf{out}}=\mathrm{softmax}\left([F(E(P_{\tau});W)]_{(d+1):(d+c),(N+1)}\right)$ be the output of the transformer for task $\tau$. We define the empirical risk over $B$ independent tasks as
@@
$L(W)=\lim_{B\to\infty}\widehat{L}(W)=-\mathbb{E}\left[\sum_{k=1}^{c}(y_{\tau,\mathsf{query}})_{k}\log((\widehat{y}_{\tau,\mathsf{out}})_{k})\right],$ (16)
@@
Theorem 4.1
@@
    Optimizing training loss $L(W)$ in (16) with training prompt length $N$ via gradient descent $W^{t+1}=W^{t}-\eta\nabla L(W^{t})$, for any $t\geq 1$, we have
@@
    Denoting $W^{*}=c(\Lambda^{-1}+G)$, we have $\|G\|_{\max}=O(c/N)$.
@@
The formal statement and proof of Theorem 4.1 can be found in Appendix F. Technically, the proof of Theorem 4.1 builds on that of Theorem 3.1, but the more complicated cross terms in the Taylor expansions of the softmax functions, which are due to the nature of multi-class classification, bring new challenges to the analysis. To address these issues, we derived new bounds on the expected errors of the cross terms in Lemma F.1, F.2, which may be of independent interest to other similar problems.
@@ 4.2 In-context inference
$\mathbb{P}\left(y_{\mathsf{query}}=\textbf{e}_{k}|x_{\mathsf{query}}\right)=\frac{f(x_{\mathsf{query}}|y_{\mathsf{query}}=\textbf{e}_{k})\mathbb{P}\left(y_{\mathsf{query}}=\textbf{e}_{k}\right)}{\sum_{j=1}^{c}f(x_{\mathsf{query}}|y_{\mathsf{query}}=\textbf{e}_{j})\mathbb{P}\left(y_{\mathsf{query}}=\textbf{e}_{j}\right)}=\mathrm{softmax}(\mu^{\top}\Lambda^{-1}x_{\mathsf{query}})_{k}.$
If we test the trained transformer with parameters $\widehat{W}$ in (19) and prompt $P_{\mathsf{test}}$, by a simple calculation, we have
@@
Theorem 4.2
(Informal) Let $P_{\mathsf{test}}$ satisfy Assumption 4.2 and $y_{\mathsf{query}}\sim\mathcal{P}^{m}_{y|x_{\mathsf{query}}}(\mu,\Lambda)$. Denote $\widehat{y}_{\mathsf{query}}$ as the prediction of the trained transformer with parameter $\widehat{W}$ in (19). Then, for the inference error defined in (3), we have
@@
The formal statement and proof of Theorem 4.2 can be found in Appendix G. We can see that the convergence rate of the inference error in multi-class classification w.r.t. $N$ and $M$ is similar to that in the binary classification, except for the constant coefficient $c$. This suggests that classification tasks with more classes may have higher errors than those with fewer classes. On the other hand, for data pair $(x,y)\sim\mathcal{P}^{m}(\mu,\Lambda)$, the Bayes-optimal classifier is $\mathbb{P}(y=\textbf{e}_{k}|x)=f(x|y)\mathbb{P}(y=\textbf{e}_{k})/f(x)=\mathrm{softmax}(\mu^{\top}\Lambda^{-1}x)_{k}$, which corresponds to a softmax regression model $\mathrm{softmax}(Wx+b)$ with parameters $W=\mu^{\top}\Lambda^{-1}$ and $b=0$. When $N,M\to\infty$, the prediction of the trained transformer is Bayes-optimal, and is equivalent to the optimal softmax regressor for multi-class classification problems with distribution $\mathcal{P}^{m}(\mu,\Lambda)$. Note that different from Assumption 4.1 which states that $\mu_{\tau},x_{\tau,\mathsf{query}}$ are sampled according to some specific distributions during training, Assumption 4.2 impose strong distributional constraints on $\mu$ or $x_{\mathsf{query}}$, which shows the strong generalization ability of the trained transformer. We also discuss the consequences when Assumption 4.2 does not hold in Remark G.1, which highlights the necessity of Assumption 4.2. Moreover, even if $M\to\infty$, the distribution variation between $y_{\mathsf{query}}$ and $\widehat{y}_{\mathsf{query}}$ does not disappear unless $N\to\infty$. Thus, the ICL ability of the trained transformers for multi-class classification is limited by the finite length of training prompts. Similar behaviors have also been observed in Zhang et al. (2023a) for in-context linear regression and in Section 3.2 for in-context binary classification.
@@ 5 Experiments
In this section, we carry out experiments to verify the theoretical claims and compare the ICL performances of transformers with other machine learning algorithms. Detailed experimental settings can be found in Appendix H.
5.1 ICL performances of the single-layer transformers
(a) $c=10$
(b) $N=80$
(c) log-log axes
(d) log-log axes
We first train single-layer transformers for in-context classification of Gaussian mixtures with different numbers of classes $c$, different lengths of training prompts $N$, and test them with different test prompt lengths $M$. The results are reported in Figure 1. We can see from Figure 1 (a,b) that the inference errors decrease as $N$ and $M$ increase, and they increase as $c$ increases. In Figure 1 (c,d), we first fix the training prompt length (test prompt length) to a large number $2000$, and then vary the test prompt length (training prompt length) from 20 to 2000. The results show that, as $M$ and $N$ become sufficiently large, the inference error, which is an approximation of $\mathbb{E}[\Delta(y_{\mathsf{query}},\widehat{y}_{\mathsf{query}})]$ (see Appendix H for detailed definitions), decreases to near-zero. This indicates that the prediction of the trained transformer approaches the Bayes-optimal classifier. All these experimental results corroborate our theoretical claims.
5.2 Comparison of transformers with other machine learning algorithms
@@
Figure 2: Inference errors for the classification of three Gaussian mixtures. Colored lines correspond to different models as follows: a single-layer transformer defined in Section 4 (1-layer, sparse), a single-layer transformer with full parameters (59) (1-layer, full), a 3-layer transformer with a GPT2 architecture (3-layer, GPT2), softmax regression (softmax), SVM with linear kernel (SVM, linear), SVM with Gaussian kernel (SVM, gaussian), 1-nearest neighbor (1-NN), and 3-nearest neighbor (3-NN). All three transformers are trained with prompt length $N=100$.
Additionally, we conduct experiments comparing the ICL performances of the transformers with other machine learning algorithms for the classification of three Gaussian mixtures. Again, the detailed experimental setting can be found in Appendix H. Form Figure 2, we can see that, when the prompt length is less than 100, all three transformer models have comparable or better performances than the traditional methods (softmax regression, SVM, $K$-nearest neighbor), which demonstrates the strong ICL capacities of transformers. When the prompt length is larger than 100, a 3-layer transformer with the GPT-2 architecture shows a decline in performance, probably because our transformer models were only trained with a small prompt length of $N=100$. Similar declined performance when the training prompt length is smaller than the test prompt length has also been observed for in-context linear regression tasks; see e.g. Zhang et al. (2023a). However, even when the prompt length is larger than 100, the single-layer transformer defined in Section 4 (1-layer, sparse) and the single-layer transformer with full parameters (59) (1-layer, full) still significantly outperform the traditional methods. This also indicates that, though the model (1-layer, sparse) we studied in this paper is relatively simple, this model is already sufficiently complex for in-context classification of Gaussian mixtures, both in theory and practice.
@@ 6 Conclusion
We studied the learning dynamics of transformers for in-context classification of Gaussian mixtures, and showed that with properly distributed data, a single-layer transformer trained via gradient descent converges to its global minimum. Moreover, we established the upper bounds of the inference errors of the trained transformers and discussed how the training and test prompt lengths influence the performance of the model. Experimental results also corroborated the theoretical claims. There are some directions worth further exploring. One potential avenue is to investigate whether the assumptions regrading the training and test prompts can be relaxed. Additionally, we have only examined single-layer transformers with linear attention and sparse parameters. The learning dynamics of multi-layer transformers with nonlinear attention (e.g., softmax) for in-context classification problems remain an interesting area for future investigation.
@@
The Appendix is organized as follows. In Section A, we provide a literature review of the related works that studied the ICL abilities of transformers. In Section B, we introduce the additional notations for the proofs in the Appendix. In Section C, we introduce some useful Lemmas we adopt from previous literature. In Sections D, E, F, G, we present the proofs of Theorem 3.1, 3.2, 4.1, 4.2 respectively. In Section H, we provide additional details of our experiments.
@@ Appendix A Related work
It has been observed that transformer-based models have impressive ICL abilities in natural language processing [Brown et al., 2020, Nye et al., 2021, Wei et al., 2022, Dasgupta et al., 2022, Zhang et al., 2022]. Garg et al. [2022] first initiated the study of the ICL abilities of transformers in a mathematical framework and they empirically showed that transformers can in-context learn linear regression, two-layer ReLU networks, and decision trees. Subsequently, numerous works have been developed to explain the ICL capacities of transformers in solving in-context mathematical problems. These works mainly use two approaches: constructing specific transformers capable of performing certain in-context learning tasks, and studying the training dynamics of transformers for such tasks.
@@
Constructions of transformers. Akyürek et al. [2022], Von Oswald et al. [2023] showed by construction that multi-layer transformers can be viewed as multiple steps of gradient descent for linear regression. Akyürek et al. [2022] also showed that constructed transformers can implement closed-form ridge regression. Guo et al. [2023] showed that constructed transformers can perform in-context learning with representations. Bai et al. [2024] proved that constructed transformers can perform various statistical machine learning algorithms through in-context gradient descent and showed that constructed transformers can perform in-context model selection. Lin et al. [2023] demonstrated that constructed transformers can approximate several in-context reinforcement learning algorithms. Fu et al. [2023], Giannou et al. [2024] further proved that constructed transformers can perform higher-order optimization algorithms like Newton’s method. Pathak et al. [2023] showed that transformers can learn mixtures of linear regressions. Giannou et al. [2023] proved that looped transformers that can emulate various in-context learning algorithms. Cheng et al. [2023] showed that transformers can perform functional gradient descent for learning non-linear functions in context. Zhang et al. [2024] showed that a linear attention layer followed by a linear layer can learn and encode a mean signal vector for in-context linear regression.
@@
Training dynamics of transformers. Mahankali et al. [2023], Ahn et al. [2024] proved that the global minimizer of the in-context learning loss of linear transformer can be equivalently viewed as one-step preconditioned gradient descent for linear regression. Zhang et al. [2023a] proved the convergence of gradient flow on a single-layer linear transformer and discussed how training and test prompt length will influence the prediction error of transformers for linear regression. Huang et al. [2023] proved the convergence of gradient descent on a single-layer transformer with softmax attention with certain orthogonality assumptions on the data features. Li et al. [2023b] showed that trained transformers can learn topic structure. Wu et al. [2023] analyzed the task complexity bound for pretraining single-layer linear transformers on in-context linear regression tasks. Tarzanagh et al. [2023] built the connections between single-layer transformers and support vector machines (SVMs). Nichani et al. [2024] showed that transformers trained via gradient descent can learn causal structure. Chen et al. [2024] proved the convergence of gradient flow on a multi-head softmax attention model for in-context multi-task linear regression. Kim and Suzuki [2024], Yang et al. [2024] proved that trained transformers can learn nonlinear features in context.
Recently, Li et al. [2024] studied the training dynamics of a single layer transformer for in-context classification problems. However, they only studied the binary classification tasks with finite patterns. They generated their data as $x=\mu_{j}+\kappa v_{k}$, where $\{\mu_{j}\}_{j=1}^{M_{1}}$ are in-domain-relevant patterns and $\{\nu_{k}\}_{k=1}^{M_{2}}$ are in-domain-irrelevant patterns, $M_{1}\geq M_{2}$ and these patterns are all pairwise orthogonal. Thus, the possible distribution of their data is finite and highly limited. In contrast, our work explores the ICL capabilities of transformers for both binary and multi-class classification of Gaussian mixtures. Specifically, our data is drawn according to $\mathcal{P}^{b}(\mu_{0},\mu_{1},\Lambda)$ or $\mathcal{P}^{m}(\mu,\Lambda)$, and the range and possible distributions of our data are infinite. Furthermore, the transformer architectures analyzed in their work also differ from those in our study, thereby highlighting the distinct contributions and independent interests of our work.
@@
Some works also studied the ICL from other perspectives. To name a few, Xie et al. [2021] explained the ICL as implicit Bayesian inference; Wang et al. [2023] explained the LLMs as latent variable models; Zhang et al. [2023b] explained the ICL abilities of transformers as implicitly implementing a Bayesian model averaging algorithm; and Li et al. [2023a] studied the generalization and stability of the ICL abilities of transformers.

### v3

@@ 4 In-context multi-class classification
$E=E(P)=\begin{pmatrix}x_{1}&x_{2}&\cdots&x_{N}&x_{\mathsf{query}}\\ y_{1}&y_{2}&\cdots&y_{N}&0_{c}\end{pmatrix}.$ (13)
@@
$\widehat{y}_{\mathsf{out}}=$ $\mathrm{softmax}\left([F(E(P);W)]_{(d+1):(d+c),(N+1)}\right)$
$=$ $\mathrm{softmax}\left(\left(\frac{1}{N}\sum_{i=1}^{N}y_{i}x_{i}^{\top}\right)Wx_{\mathsf{query}}\right).$
@@ 4.1 Training procedure
Assumption 4.2.
For each learning task $\tau\in[B]$, we assume:
@@
    $\{x_{\tau,i},y_{\tau,i}\}_{i=1}^{N}$ and $\{x_{\tau,\mathsf{query}},y_{\tau,\mathsf{query}}\}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{P}^{m}(\mu_{\tau}=(\mu_{\tau,1},\dots,\mu_{\tau,c}),\Lambda)$.
@@
    $\mu_{\tau,1}$ is sampled from $\mathsf{N}(0,I_{d})$, $\mu_{\tau,k}=U_{\tau,k,\Lambda}\mu_{\tau,1},k=2,3,\dots,c$, where $U_{\tau,k,\Lambda}=\Lambda^{1/2}U_{\tau,k}\Lambda^{-1/2}$, and $U_{\tau,k}$ are uniformly distributed over the closed set of real unitary matrices such that $U_{\tau,k}U_{\tau,k}^{\top}=I_{d}$.
We denote the distribution of $\mu_{\tau}$ as $\mathcal{P}^{m}_{\Omega}(\Lambda)$. Note that $U_{\tau,k,\Lambda}=\Lambda^{1/2}U_{\tau,k}\Lambda^{-1/2}$ can be viewed as linear transformation that preserves the inner product of vectors in the $\Lambda^{-1}$ weighted norm, and we have $\mu_{\tau,i}^{\top}\Lambda^{-1}\mu_{\tau,i}=\mu_{\tau,j}^{\top}\Lambda^{-1}\mu_{\tau,j}$, for $i,j\in[c]$. Let $\widehat{y}_{\tau,\mathsf{out}}=\mathrm{softmax}\left([F(E(P_{\tau});W)]_{(d+1):(d+c),(N+1)}\right)$ be the output of the transformer for task $\tau$. We define the empirical risk over $B$ independent tasks as
@@
$L(W)=$ $\lim_{B\to\infty}\widehat{L}(W)$
$=$ $-\mathbb{E}\left[\sum_{k=1}^{c}(y_{\tau,\mathsf{query}})_{k}\log((\widehat{y}_{\tau,\mathsf{out}})_{k})\right],$ (16)
@@
Applying gradient descent over the expected training loss in Equation (16), we have the following theorem.
Theorem 4.3.
@@
    Optimizing training loss $L(W)$ in Equation (16) with training prompt length $N$ via gradient descent $W^{t+1}=W^{t}-\eta\nabla L(W^{t})$, for any $t\geq 1$, we have
@@
    $\alpha\leq\lambda_{\min}(\nabla^{2}L(W))\leq\lambda_{\max}(\nabla^{2}L(W))\leq l,$ (18)
    where $\alpha>0$, $l<\infty$, $W\in R_{W},R_{W}=\{W\in\mathbb{R}^{d\times d}\,|\,\|W-W^{*}\|_{F}\leq\|W^{0}-W^{*}\|_{F}\}$.
@@
    Defining $G=W^{*}/c-\Lambda^{-1}$, we have $W^{*}=c(\Lambda^{-1}+G)$ and $\|G\|_{\max}=O(c/N)$.
@@
The formal statement and complete proof of Theorem 4.3 can be found in Appendix F. Technically, the proof of Theorem 4.3 builds on that of Theorem 3.3, but the more complicated cross terms in the Taylor expansions of the softmax functions, which are due to the nature of multi-class classification, bring new challenges to the analysis. To address these issues, we derive new bounds on the expected errors of the cross terms in Lemma F.1, F.2, which may be of independent interest to other similar problems.
@@ 4.2 In-context inference
$\mathbb{P}\left(y_{\mathsf{query}}=\textbf{e}_{k}|x_{\mathsf{query}}\right)$
$=$ $\frac{f(x_{\mathsf{query}}|y_{\mathsf{query}}=\textbf{e}_{k})\mathbb{P}\left(y_{\mathsf{query}}=\textbf{e}_{k}\right)}{\sum_{j=1}^{c}f(x_{\mathsf{query}}|y_{\mathsf{query}}=\textbf{e}_{j})\mathbb{P}\left(y_{\mathsf{query}}=\textbf{e}_{j}\right)}$
$=$ $\mathrm{softmax}(\mu^{\top}\Lambda^{-1}x_{\mathsf{query}})_{k}.$
If we test the trained transformer with parameters $\widehat{W}$ in Equation (19) and prompt $P_{\mathsf{test}}$, by a simple calculation, we have
@@
(a) $c=10$
(b) $N=80$
Figure 1: ’1-layer’: single-layer transformer defined in Section 4, ’3-layer’: 3-layer transformers with softmax attention. $N$: training prompt length. $c$: number of Gaussian mixtures.
Theorem 4.5.
(Informal) Let $P_{\mathsf{test}}$ satisfy Assumption 4.4 and $y_{\mathsf{query}}\sim\mathcal{P}^{m}_{y|x_{\mathsf{query}}}(\mu,\Lambda)$. Denote $\widehat{y}_{\mathsf{query}}$ as the prediction of the trained transformer with parameter $\widehat{W}$ in Equation (19). Then, for the inference error defined in Equation (3), we have
@@
The formal statement and proof of Theorem 4.5 can be found in Appendix G. We can see that the convergence rate of the inference error in multi-class classification w.r.t. $N$ and $M$ is similar to that in the binary classification, except for the constant coefficient $c$. This suggests that classification tasks with more classes may have higher errors than those with fewer classes. On the other hand, for data pair $(x,y)\sim\mathcal{P}^{m}(\mu,\Lambda)$, the distribution of $y$, $\mathbb{P}\left(y=\textbf{e}_{k}|x\right)=\mathrm{softmax}(\mu^{\top}\Lambda^{-1}x)_{k}$, can be characterized by a softmax regression model $\mathrm{softmax}(Wx+b)$ with parameters $W=\mu^{\top}\Lambda^{-1}$ and $b=0$. When $N,M\to\infty$, the prediction of the trained transformer is equivalent to the optimal softmax regressor for multi-class classification problems with distribution $\mathcal{P}^{m}(\mu,\Lambda)$. Note that different from Assumption 4.2 which states that $\mu_{\tau},x_{\tau,\mathsf{query}}$ are sampled according to some specific distributions during training, Assumption 4.4 does not impose strong distributional constraints on $\mu$ or $x_{\mathsf{query}}$, which shows the strong generalization ability of the trained transformer. We also discuss the consequences when Assumption 4.4 does not hold in Remark G.2, which highlights the necessity of Assumption 4.4. Moreover, even if $M\to\infty$, the distribution variation between $y_{\mathsf{query}}$ and $\widehat{y}_{\mathsf{query}}$ does not disappear unless $N\to\infty$. Thus, the ICL ability of the trained transformers for multi-class classification is limited by the finite length of training prompts. Similar behavior has also been observed in Zhang et al. (2023a) for in-context linear regression and in Section 3.2 for in-context binary classification.
(a) Norm
(b) Covariance
Figure 2: We generate $\Lambda_{i}=\text{diag}(\lambda_{i1},\dots,\lambda_{id})$, $i\in\{0,1,2,3\}$, where $\lambda_{ij}=|\hat{\lambda}_{ij}|$ and $\hat{\lambda}_{ij}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathsf{N}(3,1)$. All models are trained with prompt length $N=100$, tested with prompts satisfying Assumption 4.4 with $\Lambda_{0}$. $c=3$. (a) ‘same norm’: pre-training data are sampled according to Assumption 4.2 with $\Lambda_{0}$. ‘different norms’: For each $\tau$, with probability $\mathbb{P}\left(k=j\right)=1/10,\mu_{\tau,i}\sim\mathsf{N}(k,I_{d}),j=0,1,\ldots,9$. (b) ‘same covariance’: pre-training data are sampled according to Assumption 4.2 for the fixed $\Lambda_{0}$. ‘different covariances’: Sample additional $\Lambda_{1},\Lambda_{2},\Lambda_{3}$. Then, generate pre-training data according to Assumption 4.2 with $\Lambda_{0},\Lambda_{1},\Lambda_{2},\Lambda_{3}$.
@@ 5 Experiments
In this section, we report the experiment results on multi-layer, nonlinear transformers to investigate their similarities and differences to the single-layer, linear transformer we theoretically analyzed in the pervious sections. Detailed experimental settings and additional results can be found in Appendix H.
We train single-layer and multi-layer transformers for in-context classification of Gaussian mixtures with different numbers of Gaussian mixtures $c$, different lengths of training prompts $N$, and test them with different test prompt lengths $M$. The results are reported in Figure 1. We can see that for both single-layer and multi-layer transformers, the inference errors decrease as $N$ and $M$ increase, and they increase as $c$ increases, which not only verifies our theoretical claims but also shows that, the simplified model we have studied indeed exhibits behavioral similarities to the more complex multi-layer, nonlinear transformers, and some of our observations for this simplified model also hold for more complex transformers.
5.1 Varying covariances and norms
Note that in Assumption 3.2, 4.2, 3.5, 4.4, we assume that the covariance $\Lambda$ during pre-training and during inference are the same, and the means of all Gaussian components $\{\mu_{\tau,i},i\in[c]\}$ have the same $\Lambda^{-1}$ weighted norm. In Remark 3.7, G.2, we also discuss the situation when Assumptions 3.5, 4.4 do not hold and show the necessities of them. In this subsection, we consider training transformers with data of varying covariances $\Lambda$ and with Gaussian component means of unequal $\Lambda^{-1}$ weighted norms, and examine how these factors affect the ICL abilities of transformers. Results are shown in Figure 2.
From Figure 2 (a), we can see that both models perform better when their $\mu_{\tau,i}$ have the same $\Lambda^{-1}$ weighted norm (‘same norm’), however, in the ‘different norms’ setting, the performance of ‘1-layer’ deteriorates more significantly, while transformers with a more complex structure (’3-layer’) show better robustness under this distribution shift. Similar situations also happen in Figure 2 (b), where ‘3-layer’ also shows better tolerance to the covariance shifts than ‘1-layer’.
Experimental results in Figure 2 show the necessities of Assumptions 3.2, 4.2, 3.5, 4.4 for the single-layer transformers considered in this paper, and also demonstrate the better robustness of multi-layer, nonlinear transformers. Developing a better understanding of the robustness of more complex transformers is an intriguing direction for future research.
5.2 Comparison of transformers with other ML algorithms
Additionally, we conduct experiments comparing the ICL performances of the transformers with other machine learning algorithms for the classification of three Gaussian mixtures. From Figure 3, we can see that all three transformer models significantly outperform the classical methods (softmax regression, linear discriminant analysis), demonstrating the strong ICL capacities of transformers.
@@ 5.2 Comparison of transformers with other ML algorithms
Figure 3: ‘1-layer, sparse’: single-layer transformer defined in Section 4, ‘1-layer, full’: single-layer transformer with full parameters (59), ‘3-layer’: a 3-layer transformer with softmax attention, ‘softmax’: softmax regression, ‘LDA’: linear discriminant analysis. All three transformers are trained with prompt length $N=100$.
@@ 6 Conclusion
We studied the learning dynamics of transformers for in-context classification of Gaussian mixtures, and showed that with properly distributed data, a single-layer transformer trained via gradient descent converges to its global minimum. Moreover, we established the upper bounds of the inference errors of the trained transformers and discussed how the training and test prompt lengths influence the performance of the model. Experimental results also corroborated the theoretical claims. There are some directions worth further exploring. One potential avenue is to investigate whether the assumptions regarding the training and test prompts can be relaxed. Additionally, we have only examined single-layer transformers with linear attention and sparse parameters. The learning dynamics of multi-layer transformers with nonlinear attention (e.g., softmax) for in-context classification problems remain an interesting area for future investigation.
Acknowledgements
The work of Wei Shen and Cong Shen was supported in part by the US National Science Foundation (NSF) under awards ECCS-2033671, ECCS-2143559, CPS-2313110, and CNS-2002902. The work of Ruida Zhou was supported in part by NSF grants 2139304, 2146838 and the Army Research Laboratory grant under Cooperative Agreement W911NF-17-2-0196. The work of Jing Yang was supported in part by NSF award ECCS-2531023.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Acknowledgements
The Appendix is organized as follows. In Section A, we provide a literature review of the related works that studied the ICL abilities of transformers. In Section B, we introduce the additional notations for the proofs in the Appendix. In Section C, we introduce some useful Lemmas we adopt from previous literature. In Sections D, E, F, G, we present the proofs of Theorem 3.3, 3.6, 4.3, 4.5 respectively. In Section H, we provide additional results and details of our experiments.
@@ Appendix A Related work
It has been observed that transformer-based models have impressive ICL abilities in natural language processing (Brown et al., 2020; Nye et al., 2021; Wei et al., 2022; Dasgupta et al., 2022; Zhang et al., 2022). Garg et al. (2022) first initiated the study of the ICL abilities of transformers in a mathematical framework and they empirically showed that transformers can in-context learn linear regression, two-layer ReLU networks, and decision trees. Subsequently, numerous works have been developed to explain the ICL capacities of transformers in solving in-context mathematical problems. These works mainly use two approaches: constructing specific transformers capable of performing certain in-context learning tasks, and studying the training dynamics of transformers for such tasks.
@@
Constructions of transformers. Akyürek et al. (2022); Von Oswald et al. (2023) showed by construction that multi-layer transformers can be viewed as multiple steps of gradient descent for linear regression. Akyürek et al. (2022) also showed that constructed transformers can implement closed-form ridge regression. Guo et al. (2023) showed that constructed transformers can perform in-context learning with representations. Bai et al. (2024) proved that constructed transformers can perform various statistical machine learning algorithms through in-context gradient descent and showed that constructed transformers can perform in-context model selection. Lin et al. (2023) demonstrated that constructed transformers can approximate several in-context reinforcement learning algorithms. Fu et al. (2023); Giannou et al. (2024) further proved that constructed transformers can perform higher-order optimization algorithms like Newton’s method. Pathak et al. (2023) showed that transformers can learn mixtures of linear regressions. Giannou et al. (2023) proved that looped transformers that can emulate various in-context learning algorithms. Cheng et al. (2023) showed that transformers can perform functional gradient descent for learning non-linear functions in context. Zhang et al. (2024) showed that a linear attention layer followed by a linear layer can learn and encode a mean signal vector for in-context linear regression.
@@
Training dynamics of transformers. Mahankali et al. (2023); Ahn et al. (2024) proved that the global minimizer of the in-context learning loss of linear transformer can be equivalently viewed as one-step preconditioned gradient descent for linear regression. Zhang et al. (2023a) proved the convergence of gradient flow on a single-layer linear transformer and discussed how training and test prompt length will influence the prediction error of transformers for linear regression. Huang et al. (2023) proved the convergence of gradient descent on a single-layer transformer with softmax attention with certain orthogonality assumptions on the data features. Li et al. (2023c) showed that trained transformers can learn topic structure. Wu et al. (2023) analyzed the task complexity bound for pretraining single-layer linear transformers on in-context linear regression tasks. Tarzanagh et al. (2023) built the connections between single-layer transformers and support vector machines (SVMs). Nichani et al. (2024) showed that transformers trained via gradient descent can learn causal structure. Chen et al. (2024) proved the convergence of gradient flow on a multi-head softmax attention model for in-context multi-task linear regression. Kim & Suzuki (2024); Yang et al. (2024) proved that trained transformers can learn nonlinear features in context. Bu et al. (2024) studied the training convergence of transformer for tasks encoded with multiple cross-concept semantics. Li et al. (2024b) proved that a one-layer transformer can be trained to learn one-nearest neighbor for binary classification in context. Li et al. (2024a) studied the training dynamics of a single layer transformer for in-context classification problems. However, they only studied the binary classification tasks with finite patterns. They generated their data as $x=\mu_{j}+\kappa v_{k}$, where $\{\mu_{j}\}_{j=1}^{M_{1}}$ are in-domain-relevant patterns and $\{\nu_{k}\}_{k=1}^{M_{2}}$ are in-domain-irrelevant patterns, $M_{1}\geq M_{2}$ and these patterns are all pairwise orthogonal. Thus, the possible distribution of their data is finite and highly limited. In contrast, our work explores the ICL capabilities of transformers for both binary and multi-class classification of Gaussian mixtures. Specifically, our data is drawn according to $\mathcal{P}^{b}(\mu_{0},\mu_{1},\Lambda)$ or $\mathcal{P}^{m}(\mu,\Lambda)$, and the range and possible distributions of our data are infinite. Recently, (Frei & Vardi, 2024) also studied the the ICL of a linear transformer model for classifying Gaussian mixtures and showed that trained transformers can exhibit benign overfitting in-context. However, their analysis relies on additional assumptions, such as a sufficiently large signal-to-noise ratio. In contrast, our work does not rely on such assumptions. Additionally, compared to (Frei & Vardi, 2024), we consider a more general multi-class setting. We believe these distinctions highlight the independent contributions of our work.
@@
Some works also studied the ICL from other perspectives. To name a few, Xie et al. (2021) explained the ICL as implicit Bayesian inference; Wang et al. (2023) explained the LLMs as latent variable models; Zhang et al. (2023b) explained the ICL abilities of transformers as implicitly implementing a Bayesian model averaging algorithm; and Li et al. (2023b) studied the generalization and stability of the ICL abilities of transformers. Hahn & Goyal (2023) showed that ICL can arise through recombination of compositional structure found in linguistic data. Lin & Lee (2024) studied the dual operating modes for in-context classification of Gaussian mixtures. However, the analyses of both Hahn & Goyal (2023) and Lin & Lee (2024) are based on the idealized optimal next-token predictor for ICL tasks and did not discuss the training dynamics of transformer models.
Apart from studies focusing on the training convergence properties of transformers in in-context learning (ICL), many other works have investigated the convergence behavior of transformers when learning different tasks. For example, (Jelassi et al., 2022) and (Li et al., 2023a) theoretically analyzed how Vision Transformers learn and converge. (Wang et al., 2024) showed that transformers can provably learn sparse token selection, whereas fully-connected networks cannot. (Zhang et al., 2025) demonstrated that transformers can learn optimal variable selection in group-sparse classification. (Gao et al., 2024) studied the global convergence of large-scale transformers under gradient flow dynamics.